# Anisotropic Spin Ice on a Breathing Pyrochlore Lattice

### v1

@@ III.1 Isotropic Spin Ice, $\delta_{A}=\delta_{B}=0$
In the well-known case $\delta_{A}=\delta_{B}=0$, the ground state of each tetrahedron is six-fold degenerate. It is characterized by a total spin-charge of $\mathsf{Q}=0$ on each tetrahedron. This is the classical nearest-neighbor spin ice case. The number of ground states for classical spin ice can be estimated analogously to the estimation done by Pauling for water ice [37]; the total number of possible configurations $2^{N}$ for all spins times the fraction of allowed configurations for each tetrahedron $(\frac{6}{16})^{N/2}$ gives an estimate for the number of ground states $\Omega=(3/2)^{N/2}$. Here, $N$ is the number of spins in the system. Equivalently, we will use $L_{\alpha}$ as the number of unit cells in spatial direction $\alpha$; each unit cell containing four tetrahedra of the same sublattice (see Fig. 1(a)). Since on each tetrahedron, there are $4$ spins, and we can rewrite $\Omega=(3/2)^{8L_{x}L_{y}L_{z}}$. Starting from any ground state configuration for spin ice, any other configuration can be reached by forming a closed loop of spins aligning head to tail (or “in” and “out” alternatingly) and flipping them all. Since no anisotropies are at play here and the system possesses the full lattice symmetry, we will refer to this phase as isotropic spin ice.
While the here-assumed nearest neighbor spin ice model can explain many features and experimental signatures of spin ice, the long-range, dipolar interactions between spins are neglected. Theoretical investigations have shown that when considering long-range interaction, further ordering at very low temperatures occurs. However, direct experimental observations have been proven challenging due to the low temperatures required for these features to be visible [38, 14, 39, 15]. We will now show that in our anisotropic model, the ground state degeneracy always drops, even without considering the long-range dipolar interactions.
@@ III.2 Planar Ice, $\delta_{A}>0,\delta_{B}=0$
The mapping to the square ice model can be understood as follows: for all $A$ tetrahedra, fixing a single spin also fixes the spin pair’s orientation. The term spin pair refers to the spin partner in the $x,y$-plane, so $\{S_{1}^{z},S_{2}^{z}\}$ and $\{S_{0}^{z},S_{3}^{z}\}$. The opposite spin pair is, therefore, completely independent. However, the spin pair of one $A$-tetrahedron is connected via a $B$ tetrahedron to the opposite spin pair on another $A$ tetrahedron. In our numbering, if we consider a spin pair $\{S_{1}^{z},S_{2}^{z}\}$, then this is connected via a $B$ tetrahedron to the spins $\{S_{0}^{z},S_{3}^{z}\}$ of the $A$ tetrahedron above. So, we consider all $B$ tetrahedra with the same $z$ coordinate as a decoupled plane in the system. This can be seen equivalently as all $\{S_{1}^{z},S_{2}^{z}\}$ of $A$ tetrahedra with common $z$ coordinates and all $\{S_{0}^{z},S_{3}^{z}\}$ of the $A$ tetrahedron above ($z+1$), constituting a plane. Now we see that each $B$-tetrahedron in the plane acts as a vertex, the connecting $\{S_{0}^{z},S_{3}^{z}\}$ in the upper plane as the vertical links and the $\{S_{1}^{z},S_{2}^{z}\}$ in the lower plane as the horizontal links of a square lattice. An example of an $x,y$-plane of tetrahedra is shown in Fig. 2(a). In the planar ice case, individual planes of $B$ tetrahedra decouple – these are the blue tetrahedra in the figure. The lavender bonds represent the bonds of the $A$ tetrahedra, forming a square lattice and fixing the spins of the connecting $B$ tetrahedra that now act as vertices when projected onto the $x,y$-plane.
@@
It is also possible to make the same Pauling estimate for this case, where we now consider that there are $\left(\frac{12}{2^{7}}\right)^{N/4}$ ground state configurations on a $A-B$ tetrahedron pair. This gives an estimate for the ground state degeneracy of $\Omega=(3/2)^{N/4}$, which is slightly smaller than the exact value. In a single plane, flipping a closed loop of spins aligned head to tail will, again, generate all possible ground state configurations. Importantly, on the $A$ tetrahedra, however, the spin pairs must always be flipped together; only on the $B$ tetrahedra we can a choose in spin-partner be freely made. This corresponds exactly to the loop update in the planar six-vertex model.
@@
In summary, we have a system where the planes decouple in the ground state limit, and in each plane, the same constraints appear as in the six-vertex model. The ground state degeneracy, however, still grows exponentially with system size.
@@ III.3 Omni-Plane, $\delta_{A}<0,\delta_{B}=0$
In the case $\delta_{A}<0$ and $\delta_{B}=0$, we also distinguish between the $A$ and $B$ sublattice. On the $A$ sublattice, the total charge must be $\mathsf{Q}_{A}=0$ and the partial charge $\mathsf{q}_{A}^{03}=-\mathsf{q}_{A}^{12}=\pm 2$, while on the $B$ sublattice only the total charge must be $\mathsf{Q}_{B}=0$; the value of the partial charge is not fixed. This means that 2 ground states are possible on the $A$-sublattice tetrahedra, the ones with magnetization vector parallel or antiparallel with the global $z$ axis, while on the $B$-sublattice tetrahedra, all six ground states are allowed. Starting from a ground state configuration, we can now generate all possible ground states by flipping all spins on $A$ tetrahedra in a common plane since flipping a plane of $A$ tetrahedra means also flipping two spins connected by a $B$ tetrahedron, which ensures we still stay in a ground state configuration. An example of such a plane is shown in Fig. 2(a). Since any plane of connected $A$ tetrahedra could be flipped, we will refer to this phase as omni-plane phase. In each spatial direction $\alpha\in\{x,y,z\}$ there are $2L_{\alpha}$ individual planes of $A$ tetrahedra. In each plane, we can flip up to $2L_{\alpha}$ planes at the same time. We can choose $\binom{2L_{\alpha}}{i}$ possible ways to flip $i$ planes of $A$ tetrahedra. If we count all the possible flips, we will, however, count six states twice. To account for this double counting, we do not consider the initial and fully flipped state in all countings. So in total, there are $\sum_{\alpha=x,y,z}\sum_{i=1}^{2L_{\alpha}-1}\binom{2L_{\alpha}}{i}=\sum_{\alpha=x,y,z}(2^{2L_{\alpha}}-2)$ total ground states for this system.
A plane perpendicular to the $x$ or the $y$ direction is flippable only if the $A$ tetrahedra in this plane are ferromagnetically aligned; all $A$ tetrahedra in this plane must be in the same configuration. If this was not the case, flipping the plane would create a double monopole in adjacent $B$ tetrahedra. On the other hand, a plane perpendicular to the $z$ direction is only flippable if the $A$ tetrahedra in this plane are antiferromagnetically aligned. A flip in any plane creates a line of defects in the two perpendicular planes.
In summary, we have a system where the intersecting planes partially decouple, and the ground state degeneracy grows exponentially with the linear system size or the number of planes in a spatial direction.
@@ III.4 Line-Order, $\delta_{A},\delta_{B}>0$
If we have a lattice of size $\{L_{x},L_{y},L_{z}\}$, with $L_{\alpha}$ being the number of unit cells in direction $\alpha$, then the number of decoupled spin lines is given by the greatest common factor (GCF) of $L_{x}$ and $L_{y}$ multiplied by four times the height $L_{z}$ of the system. The first factor, $\text{GCF}(L_{x},L_{y})$, counting the number of spin-lines in a single $x,y$-plane, stems from the periodic boundary conditions, connecting spin-lines in a specific way. If we were to work with open boundary conditions, the number of spin lines in a single plane would be significantly higher. The factor $4$ is needed because in a single unit cell, we can find four decoupled spin-lines. So finally, we note that each spin-line can have two configurations independently, which gives a ground state degeneracy of $2^{4\cdot\text{GCF}(L_{x},L_{y})\cdot L_{z}}$.
In summary, we have a system with a ground state degeneracy, which grows with the cross-section (perpendicular to the spin-pair plane) of the system or the number of spin lines parallel to the $x,y$-plane. Along the lines of spins that lie in the strain plane, the spins are aligned head to tail.
@@ III.5 Plane-Paramagnetic, $\delta_{A}>0,\delta_{B}<0$
In the case $\delta_{A}>0,\delta_{B}<0$, we assume different signs for the two different sublattices. On both sublattices the total charge must be $\mathsf{Q}_{A/B}=0$; on the $A$ sublattice, the partial charge $\mathsf{q}_{A}^{03/12}=0$, while on the $B$ sublattice the partial charge $\mathsf{q}_{B}^{03}=-\mathsf{q}_{B}^{12}=\pm 2$. Here, 4 ground states are possible on the $A$-sublattice tetrahedra, and on the $B$-sublattice tetrahedra, 2 are allowed. The magnetization vector of the $B$ tetrahedra points along the global $z$ direction. If one spin is fixed on a $B$-sublattice tetrahedron, all spins on the whole tetrahedron are fixed, while on the $A$-sublattice, fixing one spin fixes only the opposite spin pair. Similarly to before, if we fix a single spin, we will now completely fix the neighboring $B$-tetrahedron and the opposite spin pair on the $A$-tetrahedron. Carrying this on leads to fixing a whole $x,y$-plane of $B$ tetrahedra by fixing a single spin on this plane. Since within planes, spins are ordered, but between planes, there is no correlation, we refer to this phase as plane-paramagnetic. Within a single plane, the magnetization vectors of the $B$ tetrahedra are antiferromagnetically aligned. Here, a plane is given by all $B$-tetrahedra with the same $z$ coordinates, or equivalently, all the $S_{0/3}^{z}$ of the $A$-tetrahedra above, as well as the $S_{1/2}^{z}$ of the $A$-tetrahedra below. A visual representation of such a plane is shown in Fig. 2(a), where the blue tetrahedra represent the $B$ tetrahedra and the lavender bonds the bonds on the connecting $A$ tetrahedra. Since each plane can change between the two configurations independently, the ground state degeneracy is $2^{2\cdot L_{z}}$, where $L_{z}$ is the number of unit cells in the $z$ direction. In each unit cell, there are two such planes. Naturally, all ground-state configurations can be generated by flipping planes individually.
@@
It is, however, important to note that with periodic boundary conditions, the number of tetrahedra of the same kind in $x$ and $y$ direction must be even, as otherwise, no ground state can be found. On a single line of connecting spin pairs, the pairs on the $A$ tetrahedra must be oppositely aligned, while on the $B$ tetrahedra, they must be equally aligned. So, for a system of $l=4$ $A$ tetrahedra in one direction, a possible configuration might be (+++++--++--++ - - + + - - +), where we started with an $A$ tetrahedron. The last $B$ tetrahedron closes over the boundary (++limit-from+...++ … +). For $l=3$, no such configuration can be found, as (++++--++-+ - - + + -) does not give a ground state for the last $B$ tetrahedron connected over the boundary (+limit-from+...-+ … -). Since we are working with the convention of a single unit cell already containing an even number of tetrahedra of the same kind, we will not encounter this problem.
@@
In summary, we have a system with decoupled planes, each with two possible configurations. Within a plane, neighboring $B$ tetrahedra are antiferromagnetically aligned.

### v2

@@ III.1 Isotropic Spin Ice, $\delta_{A}=\delta_{B}=0$
In the well-known case $\delta_{A}=\delta_{B}=0$, the ground state of each tetrahedron is six-fold degenerate. It is characterized by a total spin-charge of $\mathsf{Q}=0$ on each tetrahedron. This is the classical nearest-neighbor spin ice case. The number of ground states for classical spin ice can be estimated analogously to the estimation done by Pauling for water ice [49]; the total number of possible configurations $2^{N}$ for all spins times the fraction of allowed configurations for each tetrahedron $(\frac{6}{16})^{N/2}$ gives an estimate for the number of ground states $\Omega=(3/2)^{N/2}$. Here, $N$ is the number of spins in the system. Equivalently, we will use $L_{\alpha}$ as the number of unit cells in spatial direction $\alpha$; each unit cell containing four tetrahedra of the same sublattice (see Fig. 1(a)). Since on each tetrahedron, there are $4$ spins, and we can rewrite $\Omega=(3/2)^{8L_{x}L_{y}L_{z}}$. Starting from any ground state configuration for spin ice, any other configuration can be reached by forming a closed loop of spins aligning head to tail (or “in” and “out” alternatingly) and flipping them all. Since no anisotropies are at play here and the system possesses the full lattice symmetry, we will refer to this phase as isotropic spin ice. Flipping a single spin creates two monopole charges $\mathsf{Q}=\pm 1$. In our nearest neighbor model, single monopoles can move to neighboring tetrahedra by flipping a single spin (of the three equally aligned) at zero energy cost [4]. Here, the path of the monopoles, “three in - one out” or opposite, is effectively only constrained by the position of the “one” not equally aligned spin. The monopole diffusion can be considered three-dimensional since there is no actual favoring of certain alignments of spins. This picture only changes when including additional energetic considerations that lead to percolation into a fractal structure where monopoles can move [50].
While the here-assumed nearest neighbor spin ice model can explain many features and experimental signatures of spin ice, the long-range, dipolar interactions between spins are truncated beyond the nearest neighbors. Theoretical investigations have shown that when considering long-range interaction, further ordering at very low temperatures occurs. However, direct experimental observations have been proven challenging due to the low temperatures required for these features to be visible [51, 14, 52, 15]. We will now show that in our anisotropic model, the ground state degeneracy always drops, even without considering the long-range dipolar interactions.
@@ III.2 Planar Ice, $\delta_{A}>0,\delta_{B}=0$
The mapping to the square ice model can be understood as follows: for all $A$ tetrahedra, fixing a single spin also fixes the spin pair’s orientation. The term spin pair refers to the spin partner in the $x,y$-plane, so $\{S_{1}^{z},S_{2}^{z}\}$ and $\{S_{0}^{z},S_{3}^{z}\}$. The opposite spin pair is, therefore, completely independent. However, the spin pair of one $A$-tetrahedron is connected via a $B$ tetrahedron to the opposite spin pair on another $A$ tetrahedron. In our numbering, if we consider a spin pair $\{S_{1}^{z},S_{2}^{z}\}$, then this is connected via a $B$ tetrahedron to the spins $\{S_{0}^{z},S_{3}^{z}\}$ of the $A$ tetrahedron above. So, we consider all $B$ tetrahedra with the same $z$ coordinate as a decoupled plane in the system. This can be seen equivalently as all $\{S_{1}^{z},S_{2}^{z}\}$ of $A$ tetrahedra with common $z$ coordinates and all $\{S_{0}^{z},S_{3}^{z}\}$ of the $A$ tetrahedron above ($z+1$), constituting a plane. Now we see that each $B$-tetrahedron in the plane acts as a vertex, the connecting $\{S_{0}^{z},S_{3}^{z}\}$ in the upper plane as the vertical links and the $\{S_{1}^{z},S_{2}^{z}\}$ in the lower plane as the horizontal links of a square lattice. An example of an $x,y$-plane of tetrahedra is shown in Fig. 2(a). In the Fig. 2(a) an $x,y$-plane of $A$ tetrahedra is shown by the blue colored tetrahedra and the lavender bonds correspond to the $B$-tetrahedra bonds. In the planar ice case, individual planes of $B$ tetrahedra decouple, so the colors of the bonds must be inverted. The blue tetrahedra in the figure now represent the $B$ tetrahedra and the lavender bonds represent the bonds of the $A$ tetrahedra. These $A$-bonds form a square lattice and fix the spins of the connecting $B$ tetrahedra that now act as vertices when projected onto the $x,y$-plane.
@@
It is also possible to make the same Pauling estimate for this case, where we now consider that there are $\left(\frac{12}{2^{7}}\right)^{N/4}$ ground state configurations on an ABA-B tetrahedron pair. This gives an estimate for the ground state degeneracy of $\Omega=(3/2)^{N/4}$, which is slightly smaller than the exact value. In a single plane, flipping a closed loop of spins aligned head to tail will, again, generate all possible ground state configurations. Importantly, on the $A$ tetrahedra, however, the spin pairs must always be flipped together; only on the $B$ tetrahedra can we freely choose a spin partner. This corresponds exactly to the loop update in the planar six-vertex model. When we now consider a single partial charge excitation $\mathsf{q}_{A}^{03/12}$ located at an $A$ tetrahedron in a certain plane, then we can see that these are bound to stay within the same plane. The partial charge can be moved by flipping the non-spin-pairs on the $A$ tetrahedron, where the charge is located, and then creating a big loop of spins aligned head-to-tail on $B$ tetrahedra, and flipping corresponding spin-pairs on $A$ tetrahedra. Lastly, to close the loop, a non-spin-pair on an $A$ tetrahedron in the same plane must be flipped, creating a partial charge. Monopole-excitations $\mathsf{Q}_{A/B}=\pm 2$ are also bound to the two-dimensional plane they are created in and to the tetrahedral sublattice. On the $A$ tetrahedra, they are always accompanied by a partial charge. Monopole excitations can be moved to another tetrahedron of the same kind by second-order processes, crossing $A$ tetrahedra only by the corresponding spin pair.
@@
In summary, we have a system where the planes decouple in the ground state limit, and in each plane, the same constraints appear as in the six-vertex model. The ground state degeneracy, however, still grows exponentially with system size. Excitations are energetically bound to the same two-dimensional plane.
@@ III.3 Omni-Plane, $\delta_{A}<0,\delta_{B}=0$
In the case $\delta_{A}<0$ and $\delta_{B}=0$, we also distinguish between the $A$ and $B$ sublattice. On the $A$ sublattice, the total charge must be $\mathsf{Q}_{A}=0$ and the partial charge $\mathsf{q}_{A}^{03}=-\mathsf{q}_{A}^{12}=\pm 2$, while on the $B$ sublattice only the total charge must be $\mathsf{Q}_{B}=0$; the value of the partial charge is not fixed. This means that 2 ground states are possible on the $A$-sublattice tetrahedra, the ones with magnetization vector parallel or antiparallel with the global $z$ axis, while on the $B$-sublattice tetrahedra, all six ground states are allowed. Starting from a ground state configuration, we can now generate all possible ground states by flipping all spins on $A$ tetrahedra in a common plane since flipping a plane of $A$ tetrahedra means also flipping two spins connected by a $B$ tetrahedron, which ensures we still stay in a ground state configuration. An example of such a plane is shown in Fig. 2(a). Since any plane of connected $A$ tetrahedra could be flipped, we will refer to this phase as omni-plane phase. In each spatial direction $\alpha\in\{x,y,z\}$ there are $2L_{\alpha}$ individual planes of $A$ tetrahedra. Along each spatial direction, we can flip up to $2L_{\alpha}$ planes at the same time. We can choose $\binom{2L_{\alpha}}{i}$ possible ways to flip $i$ planes of $A$ tetrahedra. If we count all the possible flips, we will, however, count six states twice. To account for this double-counting, we do not consider the initial and fully flipped state in all countings. So in total, there are $\sum_{\alpha=x,y,z}\sum_{i=1}^{2L_{\alpha}-1}\binom{2L_{\alpha}}{i}=\sum_{\alpha=x,y,z}(2^{2L_{\alpha}}-2)$ total ground states for this system. A more detailed description of the double counting and the projection of the ground states is given in App. B.
A plane perpendicular to the $x$ or the $y$ direction is flippable only if the $A$ tetrahedra in this plane are ferromagnetically aligned; all $A$ tetrahedra in this plane must be in the same configuration. If this were not the case, flipping the plane would create a double monopole in adjacent $B$ tetrahedra. On the other hand, a plane perpendicular to the $z$ direction is only flippable if the $A$ tetrahedra in this plane are antiferromagnetically aligned. A flip in any plane creates a line of defects in the two perpendicular planes. Partial charge excitations and monopole excitations are dynamically immobile, as any crossing of an $A$ tetrahedron necessarily increases the energy of the system by $4|\delta_{A}|$.
In summary, we have a system where the intersecting planes partially decouple, and the ground state degeneracy grows exponentially with the linear system size or the number of planes in a spatial direction. Individual excitations on top of the ground state are fully immobile.
@@ III.4 Line-Order, $\delta_{A},\delta_{B}>0$
If we have a lattice of size $\{L_{x},L_{y},L_{z}\}$, with $L_{\alpha}$ being the number of unit cells in direction $\alpha$, then the number of decoupled spin lines is given by the greatest common factor (GCF) of $L_{x}$ and $L_{y}$ multiplied by four times the height $L_{z}$ of the system. The first factor, $\text{GCF}(L_{x},L_{y})$, counting the number of spin-lines in a single $x,y$-plane, stems from the periodic boundary conditions, connecting spin-lines in a specific way. If we were to work with open boundary conditions, the number of spin lines in a single plane would be significantly higher. The factor $4$ is needed because in a single unit cell, we can find four decoupled spin-lines. So finally, we note that each spin-line can have two configurations independently, which gives a ground state degeneracy of $2^{4\cdot\text{GCF}(L_{x},L_{y})\cdot L_{z}}$. We can create four partial charges at the corners of a closed loop within a tetrahedral plane. At the four corners, the loop must pass through the tetrahedra by not flipping the corresponding spin pair and creating a partial charge. The partial charges can not be moved independently, but only at the corner of this loop, acting therefore as membrane-operators [48, 54]. On the other hand, flipping a single spin creates two monopole-charges $\mathsf{Q}=\pm 2$. Since they can only appear in combination with a partial charge, the monopoles can only travel along the one-dimensional spin-pair line they are created on, resembling therefore lineons [48, 54]. Crossing a tetrahedron in any other direction creates additional partial charges and thus increases the energy.
In summary, we have a system with a ground state degeneracy, which grows with the cross-section (perpendicular to the spin-pair plane) of the system or the number of spin lines parallel to the $x,y$-plane. Along the lines of spins that lie in the strain plane, the spins are aligned head to tail. Partial charge excitations are mobile in a two-dimensional plane at the corner of membrane operators, and monopole excitations are only mobile along the one-dimensional spin-pair lines.
@@ III.5 Plane-Paramagnetic, $\delta_{A}>0,\delta_{B}<0$
In the case $\delta_{A}>0,\delta_{B}<0$, we assume different signs for the two different sublattices. On both sublattices the total charge must be $\mathsf{Q}_{A/B}=0$; on the $A$ sublattice, the partial charge $\mathsf{q}_{A}^{03/12}=0$, while on the $B$ sublattice the partial charge $\mathsf{q}_{B}^{03}=-\mathsf{q}_{B}^{12}=\pm 2$. Here, 4 ground states are possible on the $A$-sublattice tetrahedra, and on the $B$-sublattice tetrahedra, 2 are allowed. The magnetization vector of the $B$ tetrahedra points along the global $z$ direction. If one spin is fixed on a $B$-sublattice tetrahedron, all spins on the whole tetrahedron are fixed, while on the $A$-sublattice, fixing one spin fixes only the opposite spin pair. Similarly to before, if we fix a single spin, we will now completely fix the neighboring $B$-tetrahedron and the opposite spin pair on the $A$-tetrahedron. Carrying this on leads to fixing a whole $x,y$-plane of $B$ tetrahedra by fixing a single spin on this plane. Since within planes, spins are ordered, but between planes, there is no correlation, we refer to this phase as plane-paramagnetic. Within a single plane, the magnetization vectors of the $B$ tetrahedra are antiferromagnetically aligned. Here, a plane is given by all $B$-tetrahedra with the same $z$ coordinates, or equivalently, all the $S_{0/3}^{z}$ of the $A$-tetrahedra above, as well as the $S_{1/2}^{z}$ of the $A$-tetrahedra below. A visual representation of such a plane is shown in Fig. 2(a), where the blue tetrahedra represent the $B$ tetrahedra and the lavender bonds represent the bonds on the connecting $A$ tetrahedra. Since each plane can change between the two configurations independently, the ground state degeneracy is $2^{2\cdot L_{z}}$, where $L_{z}$ is the number of unit cells in the $z$ direction. In each unit cell, there are two such planes. Naturally, all ground-state configurations can be generated by flipping planes individually.
@@
It is, however, important to note that with periodic boundary conditions, the number of tetrahedra of the same kind in $x$ and $y$ direction must be even, as otherwise, no ground state can be found. On a single line of connecting spin pairs, the pairs on the $A$ tetrahedra must be oppositely aligned, while on the $B$ tetrahedra, they must be equally aligned. So, for a system of $l=4$ AA tetrahedra in one direction, a possible configuration might be (+++++--++--+), where we started with an $A$ tetrahedron. The last $B$ tetrahedron closes over the boundary (+++...+). For $l=3$, no such configuration can be found, as (++++--++-) does not give a ground state for the last $B$ tetrahedron connected over the boundary (++...-). Since we are working with the convention of a single unit cell already containing an even number of tetrahedra of the same kind, we will not encounter this problem. Partial charge and monopole excitations on top of the ground state are dynamically immobile as crossing any $B$ tetrahedron necessarily increases the energy by $4|\delta_{B}|$.
@@
In summary, we have a system with decoupled planes, each with two possible configurations. Within a plane, neighboring $B$ tetrahedra are antiferromagnetically aligned. Individual excitations on top of the ground state are immobile by energetic constraints.